# Spatial-Aware Decision-Making with Ring Attractors in Reinforcement Learning Systems

### v2

@@ 3.1.3 Uncertainty Quantification Integration
In the field of DRL, for any state-action pair $(s,a)$, the Q-value $Q(s,a)$ can be expressed as a function of the input state through a function approximation algorithm $\Phi_{\theta}(s)$ taking as input the current state $s$. This function approximation algorithm ($\Phi_{\theta}(s)$) can be expressed as the weight matrix of our function approximation algorithm transposed $\theta^{T}$ times the feature vector extracted from the input state $x(s)$: $Q(s,a)=\Phi_{\theta}(s)=\theta^{T}x(s)$ (Sutton & Barto, 2018). As stated in Section 3.1.2, the variance of the Gaussian functions, input to the ring attractor, will be given by the variance of the estimate value for that particular action $\sigma_{i}=\sigma_{a}$.
Among the various methods to compute the uncertainty of the action ($\sigma_{a}$) we have chosen to compute a posterior distribution with Bayesian linear regression (BLR). BLR acts as output layer for our neural network (NN) of choice. We choose a linear regression model because it does not compromise the efficiency of the NN, while at the same time it provides a distribution to compute the variance for the state-action pairs. The implementation is based on Azizzadenesheli et al. (2018), where a Bayesian value-based DQN model was instantiated to output an uncertainty-aware prediction for the state-action pairs. With this approach, the new $Q$ function is defined as:
@@
where $w_{a}$ are the weights from the posterior distribution of the BLR model. Eq. (9) represents the parameters of the final Bayesian linear layer. When provided with a state transition tuple $(s,a,r,s^{\prime})$, where $s$ is the current state, $a$ is the action taken, $r$ is the reward received, and $s^{\prime}$ is the next state. This tuple represents a single step of interaction between agent and environment in the RL framework. The model learns to adjust the weights $w_{a}$ of the BLR and the function approximation algorithm, i.e. neural networks (NNs) ($\Phi_{\theta}$), to align the Q values with the optimal action $a=argmax(\gamma\Phi_{\theta}(s)^{T}w_{a})$, (10).
@@
where $\gamma$ is the discount factor, $y$ is the expected Q-value, $\Phi_{\theta_{\text{target}}}$ are the features from the next state $s^{\prime}$ extracted by the function approximation algorithm $\Phi$ using the target network parameters, $\theta_{target}$ refers to the parameters of the target function approximation algorithm used for learning, and $\hat{a}$ is the predicted optimal action in the next state $s^{\prime}$. The construction of the Gaussian BLR prior distribution and the weights $i$ sample $w_{a,i}$ collected from the posterior distribution are performed through Thompson Sampling. This process allows us to incorporate uncertainty into our action-value estimates. For details on the construction of the Gaussian prior distribution and the specifics of the sampling process, we refer the reader to Azizzadenesheli et al. (2018). Both the mean $\bar{Q}(s,a)$ and the variance $\bar{\sigma}^{2}_{a}$ of Eq. (11) are calculated from a finite number of samples $I$.
$\bar{Q}(s,a)$ $=\frac{\sum^{i=I}_{i=0}{Q(s,a)_{i}}}{I}=\frac{\sum^{i=I}_{i=0}w_{a,i}^{T}\Phi_{\theta}(s_{t})}{I}$ (11)
$\bar{\sigma}^{2}_{a}$ $=\frac{\sum^{i=I}_{i=0}{\left(w_{a,i}^{T}\Phi_{\theta}(s_{t})-\mu_{a}\right)}^{2}}{I-1}$
0:  Input: State $s$, action $a$ from Action space $\mathcal{A}$
  Step 1: Compute Q-values and Uncertainty
  Compute mean $\bar{Q}(s,a)$ and variance $\sigma_{a}^{2}$ using Eq. 12
  Step 2: Generate Input Action Signals
  for each excitatory neuron $n$ do
     Generate Gaussian action signal using Eq. 5
     Generate Gaussian action signal with $\alpha_{n}$, $\alpha_{a}(a)$, $\bar{Q}(s,a)$ and $\sigma_{a}^{2}$ using Eq. 7
  end for
  Step 3: Reach Attractor State
  for timestep $t$ until $T$ do
     for each excitatory neuron $n$ do
        $\frac{dv_{n}}{dt}$ Update excitatory neurons using Eq. 6
     end for
     $\frac{du}{dt}$ Update inhibitory neuron using Eq. 7
  end for
  Step 4: Translate Neural Activity $V$ to $action$ selected from action space $\mathcal{A}$ using Eq. 8
  return  action
Algorithm 1 CTRNN Ring Attractor Action Selection
@@ 3.2 Deep Learning Ring Attractor Model
To further enhance the ring attractor’s integration into RL frameworks and agents, we provide a Deep Learning (DL) implementation.This approach improves model learning and integration with DRL agents. Our implementation offers both algorithmic improvements, by benefiting from DL training process, and software integration improvements, easing the deployment processes. Recurrent Neural Networks (RNNs) offer a practical approach for integrating ring attractors within DRL agents. Recent studies by Li et al. (2015) show that RNNs perform well in modeling sequential data and in capturing temporal dependencies for decision making. Like CTRNNs, RNNs mirror ring attractors’ temporal dynamics, with their recurrent connections and flexible architecture emulating the interconnected nature of ring attractor neurons. This allows modeling of weighted connections for both forward and recurrent hidden states, as shown in the Appendix A.3. The premises for modeling the structure of the RNN are as follows.
@@
Signal input as a forward pass. Forward connections from previous layers are arranged circularly, mimicking the ring’s spatial distribution. The attractor state encodes the task context, influenced by the current input and hidden state. A learnable time constant $\tau$, inherited from Eq. (2), controls input contributions and temporal evolution, enabling adaptive behavior and adaptable input contribution to the attractor state.
@@ 3.2.1 DRL Agent Integration
To shape circular connectivity within a RNN, the weighted connections in the input signal $V(s)$ and the hidden state or attractor state $U(v)$ are computed as follows:
$V(s)_{m,n}$ $=\frac{1}{\tau}\Phi_{\theta}(s)^{T}w^{I\xrightarrow{}H}_{m,n}=\frac{1}{\tau}\Phi_{\theta}(s)^{T}e^{\frac{d(m,n)}{\lambda}}$ (12)
@@
$U(v)_{m,n}$ $=h(v)^{T}w^{H\xrightarrow{}H}_{m,n}=h(\Phi_{\theta})^{T}e^{\frac{d(m,n)}{\lambda}}$
@@
This circular structure mimics the arrangement of excitatory neurons in the ring attractor. Eq. (12) shows the input signal to the recurrent layer $V_{m,n}$ from neuron $m$ from the previous layer in the DL agent to neuron $n$ in the RNN. The hidden state, $U_{m,n}$ mimics an attractor state, representing the recurrent connections in the RNN. The weighted RNN connections include fixed input-to-hidden connections ($w^{I\xrightarrow{}H}{m,n}$) to maintain the ring’s spatial structure, and learnable hidden-to-hidden connections ($w^{H\xrightarrow{}H}{m,n}$) to capture emerging action relationships. These depend on a parameter $\lambda$ that drives the decay over distance and distance between neurons $d(m,n)$ where $N$ is the total number of neurons for the RNN and $M$ is the count of neurons in the previous layer of the NN architecture. The function $\phi_{\theta}(s):\mathbb{R}^{S}\rightarrow\mathbb{R}^{M}$ maps the input state $s$ of the DL agent to a representation of characteristics that will be the input of the recurrent layer. Likewise, $\theta$ represents the parameters of this function (i.e., the weights and biases of the NN layers preceding the RNN layer, which extract relevant features from the input). The function $h(v):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is parameterized by learnable weights that map the information from previous forward passes to the current hidden state. The learnable parameter $\tau$ is the positive time constant responsible for the integration of signals in the ring. It defines the contribution of input states $\phi_{\theta}(s)$ to the current hidden state, imitating the attractor state, applied to NNs.
Finally, the action-value function $Q(s,a)$ is derived from the RNN layer’s output by applying the neurons activation function to the combined input $V(s)$ and hidden state information $U(s)$. The activation function of choice is a hyperbolic tangent $\tanh$, this function is symmetric around zero, leading to faster convergence and stability. However, the output range of $\tanh$ (-1 to 1) is not fully compatible with value-based methods, where the DL agents need to output action-value pairs in the range of the environment’s reward function. To address this issue and prevent saturation of the $\tanh$ activation function, we scale the action-value pairs by multiplying them by a learnable scalar $\beta$, as $Q(s,a)=\beta h_{t}=\beta\tanh((V(s)+U(v))=\beta\tanh((\frac{1}{\tau}\Phi_{\theta}(s_{t})^{T}w^{I\rightarrow H}+h_{t-1}(v)^{T}w^{H\rightarrow H}))$.
@@ 4 Experiments
This section presents the findings of our experiments that validate our proposed approach to integrate ring attractors into RL algorithms. To assess the effectiveness of our method, we conducted comparisons between multiple baseline models and action spaces. The evaluation encompasses two implementations: a CTRNN exogenous ring attractor and a DL approach where the ring attractor is modeled directly into DRL agents. In both implementations, the action-value pairs $Q(s,a)$ are evenly distributed across the ring circumference. For the exogenous model, each action is associated with a specific angle on the ring, Section 3.1.2. In the DL implementation, each neuron in the RNN corresponds to one action-value. The ring attractor serves as the output layer of the DL agent with the weights modeling the circular topology of the action space, Section 3.2.1. For both approaches, ring attractor agents are annotated with the suffix $RA$. We demonstrate that ring attractors significantly enhance action selection and speed up the learning process.
@@ 4.1 CTRNN Ring Attractor Model Performance
To evaluate our CTRNN ring attractor model from Section 3.1.2, integrated with BDQN (Azizzadenesheli et al., 2018) we performed experiments in the OpenAI Super Mario Bros environment  (Kauten, 2018) and OpenAI highway (Leurent, 2018). These benchmarks exhibit an spatially distributed complex decision-making scenario with 8 discrete different actions and a navigation-centric task with a continuous 1D circular actions space respectively. Fig. 2 shows that both ring attractor models (BDQNRA and BDQNRA-UA) consistently outperform standard BDQN. The uncertainty-aware version (BDQNRA-UA) shows the best overall performance, highlighting the benefits of combining ring attractors spatial distribution of the action space with uncertainty-aware action selection. Empirical evaluations revealed that CTRNN-based ring attractor models exhibited a mean computational overhead of 297.3% (SD = 14.2%) compared to baseline, significantly impacting runtime. To address this performance bottleneck and integrate the spatial understanding of the ring attractor into DRL, we developed a DL implementation of the ring attractor. This DL implementation is evaluated in the following subsections.
Figure 2: Learning speed comparison. Above: OpenAI Gym Super Mario Bros environment (Kauten, 2018) with discrete action space. Below: OpenAI highway (Leurent, 2018) with a continuous 1-D circular variable. The plot shows cumulative reward over 1 million frames for three models: Standard BDQN; BDQNRA with ring attractor behavior policy from Section 3.1.2, setting the action-value pair variance constant to $\sigma_{a}=\frac{\pi}{6}$, using this fix variance to enable smooth action transitions while preventing interference with opposing actions; and BDQNRA-UA with RA and Uncertainty Awareness (UA) implementing the uncertainty quantification model from  3.1.3 to feed into the variance of the action-value pairs. Displaying mean episodic returns over 10 averaged seeds.
@@ 4.2 DL Ring Attractor Model Performance
To evaluate the effectiveness of our DL-based ring attractor implementation from Section 3.2, we conducted experiments implementing DDQNRA, an extension of DDQN (van Hasselt et al., 2015). Fig. 3, shows that the DDQNRA model consistently outperforms standard DDQN across tasks for both navigation and game-like decision-making scenarios. This suggests that the ring attractor’s ability to encode spatial relationships in different action spaces contributes significantly to the agent’s learning speed. These results indicate that the integration of ring attractors into DRL architectures can lead to significant improvements in both learning speed and overall performance, especially in environments with strong spatial cues.
Figure 3: Performance comparison: DDQNRA vs standard DDQN (van Hasselt et al., 2015) in two environments. Above: OpenAI Super Mario Bros (Kauten, 2018), demonstrating adaptability to complex, game-like scenarios. Bellow: OpenAI highway (Leurent, 2018), showing learning speed in spatial navigation tasks. Displaying mean episodic returns over 10 averaged seeds.
@@ 4.3 Performance on Atari 100k Benchmark
In this results section, we provide a comprehensive analysis of the performance of our DRL model on the Atari 100k benchmark (Bellemare et al., 2012). We present comparisons with state-of-the-art models, highlighting the improvements achieved by our approach. We analyze the factors contributing to the significant performance increase mentioned in the introduction, breaking down results by game, and discuss notable patterns observed across different tasks. Table 1 presents a comparison of our ring attractor-based DRL model integrated with Efficient Zero (Ye et al., 2021), which evaluates performance in Atari games with a limited training size of 100,000 environment steps. The table includes results from top-performing algorithms SPR (Schwarzer et al., 2020) and CURL (Srinivas et al., 2020) for context. Our model demonstrates significant improvements over baseline methods, achieving an average 53% improvement, with over 100% gains in games with inherent spatial components, like Asterix and Boxing.
Table 1: Performance comparison on Atari 100k Benchmark (Bellemare et al., 2012). Benchmark performed across all environments where actions can be layout in one or more 2D action space planes (ring attractors). This is represented by the ring configuration column. The results are recorded at the end of training and averaged over 10 random seeds, 3 samples per seed. We display game score and overall mean and median human-normalized scores for each algorithm.
Game Agent: Reported Implemented
@@
These results reaffirm that the spatial encoding provided by ring attractors is especially beneficial in environments where spatial relationships between actions are key. The consistent performance improvement of different games indicates that our approach provides a general enhancement that benefits a wide range of RL tasks. In games where the improvement is less dramatic, we still see substantial increases in performance, suggesting that the benefits of the ring attractor extend beyond just spatially-oriented games. To ensure a fair comparison under identical experimental conditions, we re-implemented and evaluated both the baseline EffZero and our proposed EffZeroRA model using the same computational resources and experimental setup as employed throughout this study.
Ablation studies were conducted to isolate the impact of key components in our ring attractor models, detailed in Appendix A.2.1 and A.2.2 for both implementations respectively. For the exogenous model, we compared performance with correct and randomized action distributions in the ring. In the DL implementation, we removed the circular weight distribution to assess its importance.
5 Conclusion

### v3

@@ 3.1.3 Uncertainty Quantification Integration
In the field of DRL, for any state-action pair $(s,a)$, the Q-value $Q(s,a)$ can be expressed as a function of the input state through a function approximation algorithm $\Phi_{\theta}(s)$ taking as input the current state $s$. This function approximation algorithm ($\Phi_{\theta}(s)$) can be expressed as the weight matrix of our function approximation algorithm transposed $\theta^{T}$ times the feature vector extracted from the input state $x(s)$: $Q(s,a)=\Phi_{\theta}(s)=\theta^{T}x(s)$ (SuttonBartoBook, ). For further context on function approximation in RL see Appendix A.2. As stated in Section 3.1.2, the variance of the Gaussian functions, input to the ring attractor, will be given by the variance of the estimate value for that particular action $\sigma_{i}=\sigma_{a}$. Among the various methods to compute the uncertainty of the action ($\sigma_{a}$) we have chosen to compute a posterior distribution with Bayesian linear regression (BLR). BLR acts as output layer for our neural network (NN) of choice.
Algorithm 1 CTRNN Ring Attractor Action Selection
1:Input: State $s$, action $a$ from Action space $\mathcal{A}$.
2:Step 1: Compute Q-values and Uncertainty
3:Compute mean $\bar{Q}(s,a)$ and variance $\sigma_{a}^{2}$ using Eq. 11
4:Step 2: Generate Input Action Signals
5:for each excitatory neuron $n$ do
6:  Generate Gaussian action signal using Eq. 5
7:end for
8:Step 3: Reach Attractor State
9:for timestep $t$ until $T$ do $\triangleright$ choose T=50 empirically
10:  for each excitatory neuron $n$ do
11:   $\frac{\mathrm{d}v_{n}}{\mathrm{d}t}$ Update excitatory neurons using Eq. 6
12:  end for
13:  $\frac{\mathrm{d}u}{\mathrm{d}t}$ Update inhibitory neuron using Eq. 7
14:end for
15:Step 4: Translate Neural Activity $V$ to actionaction selected from action space $\mathcal{A}$ using Eq. 8
16:return action
We choose a linear regression model because it does not compromise the efficiency of the NN, while at the same time it provides a distribution to compute the variance for the state-action pairs. The implementation is based on BDQN , where a Bayesian value-based DQN model was instantiated to output an uncertainty-aware prediction for the state-action pairs. The new $Q$ function is defined as:
@@
where $w_{a}$ are the weights from the posterior distribution of the BLR model. Eq. (9) represents the parameters of the final Bayesian linear layer. When provided with a state transition tuple $(s,a,r,s^{\prime})$, where $s$ is the current state, $a$ is the action taken, $r$ is the reward received, and $s^{\prime}$ is the next state. This tuple represents a single step of interaction between agent and environment. The model learns to adjust the weights $w_{a}$ of the BLR and the function approximation algorithm, i.e. neural networks (NNs) ($\Phi_{\theta}$), to align the Q values with the optimal action $a=argmax(\gamma\Phi_{\theta}(s)^{T}w_{a})$,
@@
where $\gamma$ is the discount factor, $y$ is the expected Q-value, $\Phi_{\theta_{\text{target}}}$ are the features from the next state $s^{\prime}$ extracted by the function approximation algorithm $\Phi$ using the target network parameters, $\theta_{target}$ refers to the target function approximation algorithm parameters, and $\hat{a}$ is the predicted optimal action in the next state $s^{\prime}$. The construction of the Gaussian BLR prior distribution and the weights $i$ sample $w_{a,i}$ collected from the posterior distribution are performed through Thompson Sampling. This process allows us to incorporate uncertainty into our action-value estimates. For details on the construction of the Gaussian prior distribution and the specifics of the sampling process, we refer the reader to BDQN . Both the mean $\bar{Q}(s,a)$ and the variance $\bar{\sigma}^{2}_{a}$ of Eq. (11) are calculated from a finite number of samples $I$.
$\bar{Q}(s,a)=\frac{\sum^{i=I}_{i=0}{Q(s,a)_{i}}}{I}=\frac{\sum^{i=I}_{i=0}w_{a,i}^{T}\Phi_{\theta}(s_{t})}{I},\bar{\sigma}^{2}_{a}=\frac{\sum^{i=I}_{i=0}{\left(w_{a,i}^{T}\Phi_{\theta}(s_{t})-\mu_{a}\right)}^{2}}{I-1}$ (11)
@@ 3.2 Deep Learning Ring Attractor Model
To further enhance the ring attractor’s integration into RL frameworks and agents, we provide a Deep Learning (DL) implementation. This approach improves model learning and integration with DRL agents. Our implementation offers both algorithmic improvements, by benefiting from DL training process, and software integration improvements, easing the deployment processes. Recurrent Neural Networks (RNNs) offer an approach for integrating ring attractors within DRL agents. Recent studies by RNNRL show that RNNs perform well in modeling sequential data and in capturing temporal dependencies for decision-making. Like CTRNNs, RNNs mirror ring attractors’ temporal dynamics, with their recurrent connections and flexible architecture emulating their interconnected nature. Allowing modelling weighted connections for forward and recurrent hidden states, as shown in Appendix A.6. The premises for modeling RNN are as follows.
@@
Signal input as a forward pass. Forward connections from previous layers are arranged circularly, mimicking the ring’s spatial distribution. The attractor state encodes the task context, influenced by the current input and hidden state. A learnable time constant $\tau$, inherited from Eq. (2), controls input and temporal evolution, enabling adaptive behavior and elastic contribution to the attractor state.
@@ 3.2.1 DRL Agent Integration
To shape circular connectivity within a RNN, the weighted connections in the input signal $V(s)$ and the hidden state or attractor state $U(v)$ are computed as in (12). This circular structure mimics the arrangement of excitatory neurons in the ring attractor. Eq. (12) shows the input signal to the recurrent layer $V_{m,n}$ from neuron $m$ from the previous layer in the DL agent to neuron $n$ in the RNN. The hidden state, $U_{m,n}$ mimics an attractor state, representing the recurrent connections in the RNN. The weighted RNN connections include fixed input-to-hidden connections ($w^{I\xrightarrow{}H}{m,n}$) to maintain the ring’s spatial structure, and learnable hidden-to-hidden connections ($w^{H\xrightarrow{}H}{m,n}$) to capture emerging action relationships.
$V(s)_{m,n}$ $=\frac{1}{\tau}\Phi_{\theta}(s)^{T}w^{I\xrightarrow{}H}_{m,n}=\frac{1}{\tau}\Phi_{\theta}(s)^{T}e^{-\frac{d(m,n)}{\lambda}}$ (12)
@@
$U(v)_{m,n}$ $=h(v)^{T}w^{H\xrightarrow{}H}_{m,n}=h(\Phi_{\theta})^{T}e^{-\frac{d(m,n)}{\lambda}}$
@@
These depend on a parameter $\lambda$ that drives the decay over distance and distance between neurons $d(m,n)$ where $N$ is the total number of neurons for the RNN and $M$ is the count of neurons in the previous layer of the NN architecture. The function $\phi_{\theta}(s):\mathbb{R}^{S}\rightarrow\mathbb{R}^{M}$ maps the input state $s$ of the DL agent to a representation of characteristics that will be the input of the recurrent layer. Likewise, $\theta$ represents the parameters of this function (i.e., the weights and biases of the NN layers preceding the RNN layer, which extract relevant features from the input). The function $h(v):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is parameterized by learnable weights that map the information from previous forward passes to the current hidden state. The learnable parameter $\tau$ is the positive time constant responsible for the integration of signals in the ring. Here, $\tau$ acts as a learnable scaling factor that controls the balance between immediate input responsiveness and temporal integration within the ring attractor dynamics, with smaller values increasing sensitivity to current inputs and enabling faster adaptation, while larger values promote stability and smoother transitions between attractor states by integrating information over longer time horizons. It defines the contribution of input $\phi_{\theta}(s)$ to the current hidden state, imitating the attractor state, applied to NNs.
Finally, the action-value function $Q(s,a)$ is derived from the RNN layer’s output by applying the neurons activation function to the combined input $V(s)$ and hidden state information $U(s)$. The activation function of choice is a hyperbolic tangent $\tanh$, this function is symmetric around zero, leading to faster convergence and stability. However, the output range of $\tanh$ (-1 to 1) is not fully compatible with value-based methods, where the DL agents need to output action-value pairs in the range of the environment’s reward function. To address this issue and prevent saturation of the $\tanh$ activation function, we scale the action-value pairs by multiplying them by a learnable scalar $\beta$, as $Q(s,a)=\beta h_{t}=\beta\tanh\left(V(s)+U(v)\right)=\beta\tanh\left(\frac{1}{\tau}\Phi_{\theta}(s_{t})^{T}w^{I\rightarrow H}+h_{t-1}(v)^{T}w^{H\rightarrow H}\right)$. While our main focus is on value-based ring-attractor agents, we also provide an overview of methods and results for policy-gradient approaches in Appendix A.3.
@@ 4 Experiments
This section presents the findings of our experiments that validate our proposed approach to integrate ring attractors into RL algorithms. To assess the effectiveness of our method, we conducted comparisons between multiple baseline models and action spaces. The evaluation encompasses two implementations: a CTRNN exogenous ring attractor and a DL approach where the ring attractor is modeled directly into DRL agents. In both implementations, the action-value pairs $Q(s,a)$ are evenly distributed across the ring circumference. For the exogenous model, each action is associated with a specific discrete angle or continuous space, section 3.1.2. In the DL implementation, each neuron in the RNN corresponds to one action-value. The ring attractor serves as the output layer of the DL agent with the weights modeling the circular topology of the action space, Section 3.2.1. For both approaches, ring attractor agents are annotated with the suffix $RA$. We demonstrate that ring attractors significantly enhance action selection and speed up the learning process.
Figure 2: Learning speed comparison. Above: OpenAI Gym Super Mario Bros environment (gym-super-mario-bros, ) with discrete action space. Below: OpenAI highway (highway-env, ) with a continuous 1-D circular variable. The plot shows cumulative reward over 1 million frames for three models: Standard BDQN; BDQNRA with ring attractor behavior policy from Section 3.1.2, setting the action-value pair variance constant to $\sigma_{a}=\frac{\pi}{6}$, using this fixed variance to enable smooth action transitions while preventing interference with opposing actions; and BDQNRA-UA with RA and Uncertainty Awareness (UA) implementing the uncertainty quantification model from Section 3.1.3 to feed into the variance of the action-value pairs. Displaying mean episodic returns over 10 averaged seeds.
@@ 4.1 CTRNN Ring Attractor Model Performance
To evaluate our CTRNN ring attractor model from Section 3.1.2, integrated with BDQN (BDQN, ) we performed experiments in the OpenAI Super Mario Bros environment (gym-super-mario-bros, ) and OpenAI highway (highway-env, ). These benchmarks exhibit a spatially distributed complex decision-making scenario with 8 discrete different actions and a navigation-centric task with a continuous 1D circular actions space, respectively. Fig. 2 shows that both ring attractor models (BDQNRA and BDQNRA-UA) consistently outperform standard BDQN. The uncertainty-aware version (BDQNRA-UA) shows the best overall performance, highlighting the benefits of combining ring attractors spatial distribution of the action space with uncertainty-aware action selection. Empirical evaluations revealed that CTRNN-based ring attractor models exhibited a mean computational overhead of $297.3\%$ (SD = $14.2\%$) compared to the baseline, significantly impacting runtime. To address this performance bottleneck and integrate the spatial understanding of the ring attractor into DRL, we developed a DL implementation of the ring attractor. This DL implementation is evaluated in the following subsections.
@@ 4.2 DL Ring Attractor Model Performance
To evaluate the effectiveness of our DL-based ring attractor implementation from Section 3.2, we conducted experiments implementing DDQNRA, an extension of DDQN (double_DQN, ). Fig. 3 shows DDQNRA consistently outperforms standard DDQN, with the ring attractor’s spatial encoding ability contributing significantly to improved learning speed. These results demonstrate substantial performance gains, especially in environments with strong spatial and directional cues.
Figure 3: Performance comparison: DDQNRA vs standard DDQN (double_DQN, ) in two environments. Above: OpenAI Super Mario Bros (gym-super-mario-bros, ), demonstrating adaptability to complex, game-like scenarios. Below: OpenAI highway (highway-env, ), showing learning speed in spatial navigation tasks. Displaying mean episodic returns over 10 averaged seeds.
@@ 4.3 Performance on Atari 100k Benchmark
In this results section, we provide a comprehensive analysis of the performance of our DRL model on the Atari 100k benchmark (Atari100k, ). We present comparisons with state-of-the-art models, highlighting the improvements achieved by our approach. Table 1 presents a comparison of our ring attractor-based DRL model integrated with Efficient Zero (EffcientZero, ), which evaluates performance in Atari games with a limited training size of 100,000 environment steps. The table includes results from top-performing algorithms SPR (SPR, ) and CURL (CURL, ) for context. Our model demonstrates significant improvements over baseline methods, achieving a $53\%$ average improvement, with over $100\%$ gains in games with inherent spatial components, like Asterix and Boxing. Results demonstrate ring attractors provide consistent performance improvements across diverse RL tasks, with benefits extending beyond purely spatial tasks. We re-run baseline EffZero under identical cloud resources, to ensure fair comparison.
Ablation studies were conducted to isolate the impact of key components in our ring attractor models, detailed in Appendix A.4.1 and A.4.2 for both implementations respectively. For the exogenous model, we compared performance with correct and randomized action distributions in the ring. In the DL implementation, we removed the circular weight distribution to assess its importance. Additionally, Appendix A.5 reveals preserved attractor dynamics behaving as a low-pass filter in the CTRNN and DL both at pre and post-training stages.
Table 1: Performance comparison on Atari 100k Benchmark (Atari100k, ). Benchmark performed across all environments where actions can be laid out in one or more 2D action space planes (ring attractors). This is represented by the ring configuration column. Game score and overall mean and median human-normalized scores are recorded at the end of training and averaged over 10 random seeds, 3 samples per seed.
Game Agent Reported Implemented
@@
5 Conclusion and Future Work
Conclusion.